OpenStudy (freemap):

Which is the next term in the number pattern 1/10, 3/12, 7/14, 3/16 A. 21/18 B. 21/16 C. 19/18 D. 19/16 My answer is A?

OpenStudy (freemap):

I'm between A and C both den. are 18

jigglypuff314 (jigglypuff314):

are you sure the last one wasn't 13/16 ?

OpenStudy (freemap):

Are you talking about d the answer if so its not

jigglypuff314 (jigglypuff314):

1/10, 3/12, 7/14, 3/16 ^^ is this a 13/16

OpenStudy (pandagurl):

The pattern seems to be that both the number and the denominator are multiples of each other, and that the denominator does up by two. So based on that, we know it's either A or C. Now, which answer had both numbers have a common multiple?

OpenStudy (freemap):

If it makes it better to see this way this is a picture

OpenStudy (pandagurl):

19 and 18 don't have a common multiple, so that leaves A,

OpenStudy (freemap):

Thank you all so much for all your help. Much appreciated

OpenStudy (texaschic101):

oh..it is 13/16......then the answer is A 1/10 -- add 2 to the top and 2 to the bottom = 3/12 -- add 4 to the top and 2 to the bottom = 7/14 -- add 6 to the top and 2 to the bottom = 13/16 -- add 8 to the top and 2 to the bottom = 21/18

OpenStudy (freemap):

Thank you so much, I understood adding two to the bottom. I just couldn't figure out the top thanks for clarifying